OpenStudy (anonymous):

Simplify the radical expression: √36x^4y^5

OpenStudy (sburchette):

Assuming everything is under the radical, \[\sqrt{36x^4y^5}\] 36 is a perfect square, so 6 can be taken out from under the radical \[6\sqrt{x^4y^5}\] x^4 is the same as (x^2)^2 so you can take x^2 out from under the radical \[6x^2\sqrt{y^5}\] You can also take y^2 out from under the radical for the same reason and you will have one y left under the radical. \[6x^2y \sqrt{y}\]

OpenStudy (anonymous):